# Evolutionary dynamics with random payoff matrices

### v1

@@ 1 Introduction
In traditional evolutionary game theory, players—whether they are organisms, strategies, or agents— interact according to defined rules, with payoffs reflecting the benefits or costs tied to different strategies [sigmund:2010bo, hofbauer1998evolutionary, nowak:2006bo]. These payoffs are usually deterministic and derived from interactions between strategies, where more successful strategies tend to reproduce more quickly and ultimately dominate over time within a population.
However, in many real-world scenarios, environmental fluctuations and incomplete information render payoffs unpredictable  [fudenberg:1992bv, gross2009generalized, GokhalePNAS2010, Galla2013, DuongHanJMB2016, hilbe2018indirect, wang2024evolution]. For example, the payoff for a predator strategy may vary due to changes in prey abundance or the presence of competing predators. Therefore, it is of vital importance for the applicability of evolutionary game theory to integrate randomness and uncertainty into the evolutionary models. This need has given rise to the concept of evolutionary games with random payoffs.
@@
We refer the reader to [fudenberg:1992bv, gross2009generalized, GokhalePNAS2010, Galla2013, DuongHanJMB2016, hilbe2018indirect, wang2024evolution] again, as well as the recent survey articles [traulsen2023future, leimar2023game] for further information.
@@ 2 Evolutionarily stable strategies in random games
Evolutionarily stable strategy (ESS), introduced by Maynard Smith and Price in 1973 [SP73], is a foundational concept in evolutionary game theory. This concept refines the notion of Nash equilibrium (see below) and has many applications in fields such as ecology, economics and social sciences.
@@
A mixed strategy $p\in\Delta_{n}$ is an ESS if it satisfies the two conditions [SP73]:
@@
The support of an ESS $p$ is defined as $\delta(p)=\{i:p_{i}>0\}$. Biologically, the existence of a mixed ESS and the size of its support provides useful insight into the co-existence and diversity of the strategies ($|\delta(p)|=1$ corresponds to monomorphism and $|\delta(p)|>1$ to polymorphism). Thus, when the payoff entries $A_{ij}$ are random variables, two natural questions arise: what is the probability that an ESS exists, and how large is the size of the support of the ESS? In a series of works [Haigh1988, haigh1989large, Haigh1990, hart2008evolutionarily], the author addresses these questions for some specific and small values of $n$ and distributions of the payoff entries, while the paper [kontogiannis2009support] investigates the case of large $n$ and broader classes of distributions. A remarkable result obtained in [kontogiannis2009support] is that when payoff entries behave either as uniform, or normally distributed random variables, almost all ESSs have support sizes of $o(n)$, where $n$ is the number of possible types for a player. With the aforementioned biological relevance of the concept of the support of an mixed ESS, the results of these work have implications for understanding the diversity of stable strategies within a population and gives insights into the relationship between the number of strategies and the robustness of an ESS. They also highlight how randomness influences population dynamics, particularly in terms of stability, coexistence, and the prevalence of certain strategies over others. More recently, intriguing connections between these papers to random quadratic optimization problems have been established [pittel2018random, chen2021sparsity].
@@ 3 Equilibria of random replicator-mutator dynamics
The replicator-mutator equation, which is a set of ordinary differential equations, describes the evolution dynamics of different strategies in a population, where selection and mutation are both captured. It is an well-established mathematical framework that integrates the unavoidable mutation observed in various biological and social settings [traulsen:2009aa, rand2013evolution, zisis2015generosity, mcnamara2013towards, adami2016evolutionary, wang2020robust]. This framework has been utilised in many application domains, including evolution of collective behaviours [Imhof-etal2005, nowak:2006bo], social networks dynamics [Olfati2007], language evolution [Nowaketal2001], population genetics [Hadeler1981], and autocatalytic reaction networks [StadlerSchuster1992].
@@
The replicator-mutator is given by [Komarova2001JTB, Komarova2004, Komarova2010, Pais2012]
@@
The explicit formulas for the average payoffs $\{f_{i}\}_{i=1}^{n}$ depend on the type of the game that the players evolve. In symmetric games, a player’s payoff within a group does not depend on the ordering of its members, while in asymmetric games, it does. In both cases, the average payoffs are all multivariate polynomials in $x_{1},\ldots,x_{n}$, thus finding an equilibrium of a $d$-player $n$-strategy evolutionary game using the replicator-mutation dynamics (2) is equivalent to finding a root in $[0,1]^{n}$ of the system of polynomial equations (3). Accordingly, when the payoff entries are random variables, finding an equilibrium of the replicator-mutator dynamics (2) becomes finding a root in the $n$-dimensional unit cube of the system of random polynomials. This correspondence has been revealed and explored in a series of our works [DuongHanDGA2016, DuongHanJMB2016, DuongTranHanDGA, DuongTranHanJMB, duong2019persistence, can2022expected, duong2023random]. Let $\mathcal{N}$ be the expected number of internal equilibria in $d$-player two-strategy games in the replicator dynamics. It has been known [DuongHanDGA2016, DuongHanJMB2016, can2022expected], that in symmetric games, when the payoff entries are iid Gaussian random variables then
@@
Remarkably, by revealing that the random polynomial arising from asymmetric games is exactly the renowned the Kostlan-Shub-Smale random polynomial, we have recently shown that [duong2023random]
@@
for some $c>0$ depending only on $\varepsilon$. In [DuongTranHanDGA, DuongHanDGA2020] we have partially extended these results to the replicator-mutator dynamics and the case of correlated payoff entries respectively. It is also worth mentioning related works  [huang:2010aa, huang2012impact] on fixation probabilities and the average population fitness of random games capturing random mutations.
Open problems: We have conjectured, supported by numerical evidence in [duong2023random], that universality results also apply to the class of random polynomials derived from symmetric evolutionary games. Can this conjecture rigorously be proven?
@@ 4 Social dilemmas under fluctuating environments
Social dilemmas arise in many social, biological and economic contexts, including multi-national conservation of natural resources, over-fishing and overgrazing of common property [fatima2024learning]. In a simplified setting, social dilemmas are often modeled as a pairwise game with two strategies, cooperation (C) and defection (D), with a general payoff matrix given by
@@
In many scenarios, the payoff entries are not fixed and fluctuate around certain known values, which are for example estimated through data analysis or given by domain experts. Hence studying the evolutionary outcomes of social dilemmas under the fluctuating environments has been of both practical and theoretical interest. In [DuongHan2021, chen2024number], using the replicator-mutator dynamics (4), we analyze the probability distribution of the number of equilibria as a function of the mutation strength for the four aforementioned social dilemmas, see Figure 1 for an illustration. Compared with the generic random games considered in Section 3, in these works the payoff entries fluctuate around the standard deterministic values that respect the orderings of the corresponding game. One can view this approach as a randomization of the concrete games to capture uncertainty. By varying the type and amplitude of the noise, we reveal its influence on the evolutionary outcomes of the dynamics. In a similar setting, in [AmaralRoyal2020, AmaralPRE2020] the authors show a rich spectrum of possible equilibria, highlighting the relationship between heterogeneity and cooperation. From a more applied side, several experimental studies estimate the entries in the fitness matrix from in vitro data of interactions between microbial populations, for evolutionary game theory analysis [turner1999prisoner, kaznatcheev2019fibroblasts, wolfl2022contribution].
Figure 1: Probabilities of observing a certain number of equilibrium points for each social dilemma game, for different mutation strengths. Payoffs are drawn from uniform distributions. Reproduced from [DuongHanDGA2020].
@@ 5 Dynamical properties of random games
We are also interested in dynamical properties of this dynamics. As a dynamical system, it may exhibit different qualitative behaviours such as converging to a fixed point, having limit circles or becoming chaotic when $n\geq 3$ according to the Poincaré-Bendixon theorem. Together with equilibrium properties, the dynamical properties of the replicator-mutator dynamics often provide further useful insights into the underlying real-world systems. For instance, in the aforementioned applications of the replicator-mutator in language evolution and behaviour adoption in social networks [Nowaketal2001, Komarova2001JTB, Komarova2004], its limit cycles correspond to, respectively, oscillations in the dominance of the different grammars in the population, and oscillations of behavior preference, for example cycles in trends or fashions. Thus studying dynamical properties of the replicator-mutator dynamics is of both mathematical and practical interests. However, the replicator-mutator dynamics (2) is a nonlinear system with several parameters; hence, in general it is analytically intractable.
@@
In [nowak:1993pv, mitchener2004chaos], by sampling the random payoff entries and simulating the replicator-mutator dynamics, the authors demonstrate that the replicator dynamics for $n\geq 4$ can generate limit cycles and chaos for particular choices of fitness. Motivated by this prediction, in [Pais2012], the authors rigorously prove Hopf bifurcations for the replicator-mutator dynamics with $n\geq 3$ and characterize the existence of stable limit cycles for a class of circulant fitness matrices.
@@
In recent years, the hybrid approach of combining sampling simulations and analytically proving the cyclic behaviour for the evolutionary dynamics has increasingly attracted attention from many researchers. In [park2020cyclic], the authors consider the rock-paper-scissor games, which is an instance of the replicator-mutator dynamics described above with three strategies, with random payoff entries modeling the interactions occurring when a new mutant type arises. They show that the conditions needed for an interaction matrix to provide cyclic dominance are more restrictive than those for non-cyclic dominance, thus the chance to form cyclic dominance is small, explaining its rareness in nature. In a similar spirit and also for the rock-paper-scissor games, in [kleshnina2021mistakes], the authors show analytically that randomness in the payoff entries, interpreted as heterogeneity or individuals’ behavioural mistakes, may break a cyclic relationship and lead to a stable equilibrium in pure or mixed strategies. In addition, in [feng2022noise], using tools in stochastic differential equations, the authors investigate how environmental noise may affect cyclic dynamical behavior of the rock–paper–scissors game. They deduce conditions for stochastic stability to determine whether environmental noise can induce a heteroclinic cycle. The aforementioned studies reveal the evolutionary significance of environmental stochastic fluctuations in biological populations with or without cyclic dominance.
@@
Open problem: The replicator-mutator dynamics for three or more strategies also arise naturally in recent works on multi-population models for AI governance and healthcare system [alalawi2020toward, alalawi2024, bova2024both]. For instance, in the simplified model for AI governance [alalawi2024], the three populations consist of the regulators, AI users and the AI firms. Similarly, for applications in healthcare, they are the patients, the private sectors and the public sectors [alalawi2020toward]. Motivated by the aforementioned works on the rock-paper-scissors games, one interesting question is: can one derive dynamical properties such as limit cycles and bifurcations for these multi-population models? {mdframed}[backgroundcolor=green, linecolor=black,font=,frametitle=Summary of Open Problems]
@@ Acknowledgment
[…]

### v2

@@ 1 Introduction
In traditional evolutionary game theory, players—whether they are organisms, strategies, or agents— interact according to defined rules, with payoffs reflecting the benefits or costs tied to different strategies [52, 30, 44]. These payoffs are usually deterministic and derived from interactions between strategies, where more successful strategies tend to reproduce more quickly and ultimately dominate over time within a population.
However, in many real-world scenarios, environmental fluctuations and incomplete information render payoffs unpredictable  [20, 23, 22, 21, 11, 29, 57]. For example, the payoff for a predator strategy may vary due to changes in prey abundance or the presence of competing predators. Therefore, it is of vital importance for the applicability of evolutionary game theory to integrate randomness and uncertainty into the evolutionary models. This need has given rise to the concept of evolutionary games with random payoffs.
@@
We refer the reader to [20, 23, 22, 21, 11, 29, 57] again, as well as the recent survey articles [54, 40] for further information.
@@ 2 Evolutionarily stable strategies in random games
Evolutionarily stable strategy (ESS), introduced by Maynard Smith and Price in 1973 [41], is a foundational concept in evolutionary game theory. This concept refines the notion of Nash equilibrium (see below) and has many applications in fields such as ecology, economics and social sciences.
@@
A mixed strategy $p\in\Delta_{n}$ is an ESS if it satisfies the two conditions [41]:
@@
The support of an ESS $p$ is defined as $\delta(p)=\{i:p_{i}>0\}$. Biologically, the existence of a mixed ESS and the size of its support provides useful insight into the co-existence and diversity of the strategies ($|\delta(p)|=1$ corresponds to monomorphism and $|\delta(p)|>1$ to polymorphism). Thus, when the payoff entries $A_{ij}$ are random variables, two natural questions arise: what is the probability that an ESS exists, and how large is the size of the support of the ESS? In a series of works [25, 26, 27, 28], the author addresses these questions for some specific and small values of $n$ and distributions of the payoff entries, while the paper [39] investigates the case of large $n$ and broader classes of distributions. A remarkable result obtained in [39] is that when payoff entries behave either as uniform, or normally distributed random variables, almost all ESSs have support sizes of $o(n)$, where $n$ is the number of possible types for a player. With the aforementioned biological relevance of the concept of the support of an mixed ESS, the results of these work have implications for understanding the diversity of stable strategies within a population and gives insights into the relationship between the number of strategies and the robustness of an ESS. They also highlight how randomness influences population dynamics, particularly in terms of stability, coexistence, and the prevalence of certain strategies over others. More recently, intriguing connections between these papers to random quadratic optimization problems have been established [50, 10].
@@ 3 Equilibria of random replicator-mutator dynamics
The replicator-mutator equation, which is a set of ordinary differential equations, describes the evolution dynamics of different strategies in a population, where selection and mutation are both captured. It is an well-established mathematical framework that integrates the unavoidable mutation observed in various biological and social settings [55, 51, 60, 42, 1, 58]. This framework has been utilised in many application domains, including evolution of collective behaviours [33, 44], social networks dynamics [47], language evolution [45], population genetics [24], and autocatalytic reaction networks [53].
@@
The replicator-mutator is given by [38, 36, 37, 48]
@@
The explicit formulas for the average payoffs $\{f_{i}\}_{i=1}^{n}$ depend on the type of the game that the players evolve. In symmetric games, a player’s payoff within a group does not depend on the ordering of its members, while in asymmetric games, it does. In both cases, the average payoffs are all multivariate polynomials in $x_{1},\ldots,x_{n}$, thus finding an equilibrium of a $d$-player $n$-strategy evolutionary game using the replicator-mutation dynamics (2) is equivalent to finding a root in $[0,1]^{n}$ of the system of polynomial equations (3). Accordingly, when the payoff entries are random variables, finding an equilibrium of the replicator-mutator dynamics (2) becomes finding a root in the $n$-dimensional unit cube of the system of random polynomials. This correspondence has been revealed and explored in a series of our works [12, 11, 16, 17, 7, 8, 15]. Let $\mathcal{N}$ be the expected number of internal equilibria in $d$-player two-strategy games in the replicator dynamics. It has been known [12, 11, 8], that in symmetric games, when the payoff entries are iid Gaussian random variables then
@@
Remarkably, by revealing that the random polynomial arising from asymmetric games is exactly the renowned the Kostlan-Shub-Smale random polynomial, we have recently shown that [15]
@@
for some $c>0$ depending only on $\varepsilon$. In [16, 13] we have partially extended these results to the replicator-mutator dynamics and the case of correlated payoff entries respectively. It is also worth mentioning related works  [31, 32] on fixation probabilities and the average population fitness of random games capturing random mutations.
Open problems: We have conjectured, supported by numerical evidence in [15], that universality results also apply to the class of random polynomials derived from symmetric evolutionary games. Can this conjecture rigorously be proven?
@@ 4 Social dilemmas under fluctuating environments
Social dilemmas arise in many social, biological and economic contexts, including multi-national conservation of natural resources, over-fishing and overgrazing of common property [18]. In a simplified setting, social dilemmas are often modeled as a pairwise game with two strategies, cooperation (C) and defection (D), with a general payoff matrix given by
@@
In many scenarios, the payoff entries are not fixed and fluctuate around certain known values, which are for example estimated through data analysis or given by domain experts. Hence studying the evolutionary outcomes of social dilemmas under the fluctuating environments has been of both practical and theoretical interest. In [14, 9], using the replicator-mutator dynamics (4), we analyze the probability distribution of the number of equilibria as a function of the mutation strength for the four aforementioned social dilemmas, see Figure 1 for an illustration. Compared with the generic random games considered in Section 3, in these works the payoff entries fluctuate around the standard deterministic values that respect the orderings of the corresponding game. One can view this approach as a randomization of the concrete games to capture uncertainty. By varying the type and amplitude of the noise, we reveal its influence on the evolutionary outcomes of the dynamics. In a similar setting, in [4, 5] the authors show a rich spectrum of possible equilibria, highlighting the relationship between heterogeneity and cooperation. From a more applied side, several experimental studies estimate the entries in the fitness matrix from in vitro data of interactions between microbial populations, for evolutionary game theory analysis [56, 34, 59].
Figure 1: Probabilities of observing a certain number of equilibrium points for each social dilemma game, for different mutation strengths. Payoffs are drawn from uniform distributions. Reproduced from [13].
@@ 5 Dynamical properties of random games
We are also interested in dynamical properties of this dynamics. As a dynamical system, it may exhibit different qualitative behaviours such as converging to a fixed point, having limit circles or becoming chaotic when $n\geq 3$ according to the Poincaré-Bendixon theorem. Together with equilibrium properties, the dynamical properties of the replicator-mutator dynamics often provide further useful insights into the underlying real-world systems. For instance, in the aforementioned applications of the replicator-mutator in language evolution and behaviour adoption in social networks [45, 38, 36], its limit cycles correspond to, respectively, oscillations in the dominance of the different grammars in the population, and oscillations of behavior preference, for example cycles in trends or fashions. Thus studying dynamical properties of the replicator-mutator dynamics is of both mathematical and practical interests. However, the replicator-mutator dynamics (2) is a nonlinear system with several parameters; hence, in general it is analytically intractable.
@@
In [46, 43], by sampling the random payoff entries and simulating the replicator-mutator dynamics, the authors demonstrate that the replicator dynamics for $n\geq 4$ can generate limit cycles and chaos for particular choices of fitness. Motivated by this prediction, in [48], the authors rigorously prove Hopf bifurcations for the replicator-mutator dynamics with $n\geq 3$ and characterize the existence of stable limit cycles for a class of circulant fitness matrices.
@@
In recent years, the hybrid approach of combining sampling simulations and analytically proving the cyclic behaviour for the evolutionary dynamics has increasingly attracted attention from many researchers. In [49], the authors consider the rock-paper-scissor games, which is an instance of the replicator-mutator dynamics described above with three strategies, with random payoff entries modeling the interactions occurring when a new mutant type arises. They show that the conditions needed for an interaction matrix to provide cyclic dominance are more restrictive than those for non-cyclic dominance, thus the chance to form cyclic dominance is small, explaining its rareness in nature. In a similar spirit and also for the rock-paper-scissor games, in [35], the authors show analytically that randomness in the payoff entries, interpreted as heterogeneity or individuals’ behavioural mistakes, may break a cyclic relationship and lead to a stable equilibrium in pure or mixed strategies. In addition, in [19], using tools in stochastic differential equations, the authors investigate how environmental noise may affect cyclic dynamical behavior of the rock–paper–scissors game. They deduce conditions for stochastic stability to determine whether environmental noise can induce a heteroclinic cycle. The aforementioned studies reveal the evolutionary significance of environmental stochastic fluctuations in biological populations with or without cyclic dominance.
@@
Open problem: The replicator-mutator dynamics for three or more strategies also arise naturally in recent works on multi-population models for AI governance and healthcare system [3, 2, 6]. For instance, in the simplified model for AI governance [2], the three populations consist of the regulators, AI users and the AI firms. Similarly, for applications in healthcare, they are the patients, the private sectors and the public sectors [3]. Motivated by the aforementioned works on the rock-paper-scissors games, one interesting question is: can one derive dynamical properties such as limit cycles and bifurcations for these multi-population models? {mdframed}[backgroundcolor=green, linecolor=black,font=,frametitle=Summary of Open Problems]
@@ Acknowledgment
T.A.H. is supported by EPSRC (grant EP/Y00857X/1) and the Future of Life Institute. M.H.D is supported by EPSRC (grant EP/Y008561/1) and a Royal International Exchange Grant IES-R3-223047.